OpenStudy (anonymous):

Solve this problem: A girl scout troop sold cookies. The girls sold 5 more boxes the second week than they did the first. They doubled the sales of the second week for the third week to sell a total of 431 boxes of cookies. How many boxes of cookies did they sell each week?

jimthompson5910 (jim_thompson5910):

let x = amount sold in first week y = amount sold in second week z = amount sold in third week

jimthompson5910 (jim_thompson5910):

The girls sold 5 more boxes the second week than they did the first means y = x+5

jimthompson5910 (jim_thompson5910):

They doubled the sales of the second week for the third week means z = 2y

jimthompson5910 (jim_thompson5910):

they sold a total of 431 boxes, which means x+y+z = 431

jimthompson5910 (jim_thompson5910):

Now solve y = x+5 for x to get x = y - 5

jimthompson5910 (jim_thompson5910):

then plug in x = y-5 and z = 2y into x+y+z = 431 x+y+z = 431 y-5+y+2y = 431 Solve that for y. Then use that solution of y to find solutions for x and z

OpenStudy (anonymous):

So, they sold 104 boxed the first week. 109 the second and 218 the third.

jimthompson5910 (jim_thompson5910):

I'm getting the same and I just checked the answer and it all works out nice job

OpenStudy (anonymous):

Thanks.

jimthompson5910 (jim_thompson5910):

yw